ROTATIONALLY RESOLVED SPECTROSCOPY OF THE ELECTRONICALLY EXCITED C AND D STATES OF XeKr AND XeAr

LORENA PITICCO, MARTIN SCHÄFER, and \underline{FRÉDÉRIC MERKT}, ETH Zürich, Laboratorium für Physikalische Chemie, Wolfgang-Pauli-Strasse 10, 8093 Zürich, Switzerland.

Informations on excited electronic states of the heteronuclear rare-gas dimers XeRg (Rg=Kr, Ar) available in the literature are limited to the vibrational structure of several band systems in the VUV range of the electromagnetic spectrum.^{a} Using a near-Fourier-transform-limited vacuum-ultraviolet laser system^{b} spectra of the C ← X and D ← X band systems of several isotopomers of XeKr and XeAr were recorded at high resolution in the wavenumber range from 77 000 cm^{−1} to 77 350 cm^{−1} by resonance-enhanced two-photon ionization spectroscopy. The rotational and vibrational structures of the C ← X and D ← X band systems could be fully resolved and assigned on the basis of isotopic shifts, combination differences and the ground state microwave spectra^{c}. The orbital hyperfine structure of the C1 state could be resolved for the ^{129}Xe^{40}Ar and 131Xe^{40}Ar isotopomers. Potential energy functions and a full set of spectroscopic parameters were derived for the ground and excited states.

aD. M. Mao, X. K. Hu, S. S. Dimov, R. H. Lipson, J. Phys. B 29, L89 (1996). O. Zehnder, F. Merkt, Mol. Phys. 106, 1215 (2008). O. Zehnder, F. Merkt, J. Chem. Phys. 128, 014306 (2008).

bU. Hollenstein, H. Palm and F. Merkt, Rev. Sci. Instr. 71, 4023 (2000).

cW. Jäger, Y. Xu, M. C. L. Gerry, J. Chem. Phys. 99, 919 (1993).